OpenStudy (anonymous):

One gallon of paint usually covers about 350 square feet. How many square feet will 2.5 gallons cover?

OpenStudy (nubeer):

total gallon cover area = area cover by one gallon * no. of gallons = 2.5* 350

OpenStudy (anonymous):

ratio x sq ft = 2.5 gal 350 sq ft = 1 gal then x sq ft 2.5 gal ------- = ---------- 350 sq ft 1 gal then x=350(2.5)/1 = ______?

OpenStudy (anonymous):

GALLON SQUARE FEET 1 350 2.5 let this be N WHEN YOU CROSS MULTIPLY N=350*2.5=875 2.5 GALLONS WILL COVER 875 SQRE FEET